A firm offers terms of 1.4/10, net 30.

  

a.

What effective annual interest rate does the firm earn when a customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What effective annual interest rate does the firm earn if the terms are changed to 2.4/10, net 30, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What effective annual interest rate does the firm earn if the terms are changed to 1.4/10, net 45, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
d. What effective annual interest rate does the firm earn if the terms are changed to 1.4/15, net 30, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a. Effective annual interest rate = [1 + (1.4/98.6)][365 / (30 - 10)] - 1 = 0.2934 or 29.34%

b. Effective annual interest rate = [1 + (2.4/97.6)][365 / (30 - 10)] - 1 = 0.5579 or 55.79%

c. Effective annual interest rate = [1 + (1.4/98.6)][365 / (45 - 10)] - 1 = 0.1584 or 15.84%

d. Effective annual interest rate = [1 + (1.4/98.6)][365 / (30 - 15)] - 1 = 0.4093 or 40.93%